A flat retangular pice of aluminum has a perimeter of 60 inches. The length is 14 inches longer than the width. Find the width.
Answer by graphmatics(170)

Let w be the width then the length is w+14. Because the perimeter is 69 we have
2*w+2*(w+14)=60
4*w+28=60
w=32/4=8